# General Chemistry Exam with Artificial Intelligence Jan Cvengros,a and Gerd Kortemeyerb†

### v1

@@ 3.4.1 Partial-credit threshold
Experiences from earlier studies suggest that AI is more likely to grade correct answers as incorrect (i.e., produce false negatives) that the other way around (false positives) 39. One possible filter is thus to binarize the partial grading decisions and treat negatives as suspicious. For the purpose of filtering the AI grading decisions, we turn the partial-credit judgements into binary decisions based on a threshold $t$ for “correctness.” If a result is “incorrect" or “incorrect enough” based on our threshold, we discard the AI-judgement and, in a production setting, would delegate the grading of that problem part for that student to a human grader.
@@
Cases where the AI marks a part as “‘correct”’ but the TA”s score falls below $t$.
@@ 3.4.2 Risk threshold
Psychometrics originated in the early 20th century as a means to standardize the measurement of psychological traits such as intelligence and aptitude. The field introduced techniques like factor analysis to identify latent variables underlying observed responses, reducing them to common dimensions. A prominent early example was the general intelligence factor, $g$, proposed to capture overall cognitive ability 40, 41, 42. Over time, such broad constructs gave way to domain-specific abilities shaped by learning, and the focus of psychometrics gradually shifted from evaluating individuals to evaluating the assessments themselves, providing tools to ensure the validity, reliability, and fairness of tests.
@@
where $\theta_{i}$ is the latent “ability” of student $i$, and $a_{j}$ and $b_{j}$ are the estimated discrimination and difficulty of part $j$, respectively 36, 43.
The AI’s actual normalized score $s_{ij}$ is then compared to this expectation 44, and we define the “risk” of accepting that AI judgment as the absolute deviation
@@ 3.4.3 Problem types
Previous experiences have shown that multimodal AI does oftentimes succeed in reading handwriting, but still fails interpreting scientific diagrams and graphs 45, 25, 39, 46. A straightforward filtering mechanism is thus to exclude any problem parts that have graphical components, namely “Drawing” and “Graphing” (see Fig. 3), i.e., filter for the remaining problem parts that are purely textual or require nothing more than recognizing the position of a cross in multiple-choice problems..
@@ 4.1 Qualitative results
A comparison of the exam sheets graded by the TAs and score sheets provided by AI, revealed some interesting features of the AI-grading.
@@ 4.1.2 False negative
Fig.  6:  An example of a false negative (problem 6-C-b).
@@
Fig.  7:  An example of a false negative, where AI assigned only partial credit (problem 2-C-a).
@@
Fig.  8:  Incorrect interpretation of a multiple-response problem (problem 6-A).
In Fig. 8, AI does not correctly recognize the position of a cross. It claims that the second last answer contains the cross in the left column named “richtig” (correct) and deducts points. It is unclear how this may have happened, except that the student markings are not really crosses.
Fig.  9:  Incorrect interpretation of a labels (problem 4-A-a).
@@ 4.2 Unfiltered quantitative results
Fig.  10:  Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the six problems on the exam. Each datapoint represents one student.
As Fig. 10 shows, for the individual six problems, the AI-derived scores correlated with human scores at coefficients of determination ($R^{2}$) ranging from $0.61$ to $0.85$. The reliability on a per-problem base is low and not suitable for grading without additional measures, yet the aggregate total score shown in the top-left panel of Fig. 11 achieved an $R^{2}$ of $0.91$ — that this $R^{2}$ is “bigger than the sum of its parts” may initially appear surprising. The enhancement results from the partial cancellation of random scoring errors across problems: when individual deviations offset one another in the sum, overall noise diminishes relative to the true-score variance. The resulting increase in the signal-to-noise ratio produces a more reliable composite measure of student proficiency than any single item alone.
Fig.  11:  Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the full exam.
As Table 1 shows, the multiclass quality of the grading varies between question types. The first column lists the question type with additional rows across all question types and a row for combined question types. The second column shows the number of these parts, the third row the number of student-part samples, the forth column the number of classes $n$, the four following columns the quality measures, and the rightmost column the normed F1 (Equation 1. The table is sorted in order of decreasing normed F1 to highlight the questions types that are most accurately graded.
Table 1:  Quality measures for the AI grading results in order of decreasing F1.
@@
Not surprisingly for a language model, short and long textual answers emerge as the most reliably graded one. Fortunately for chemistry education, equations of chemical reactions are also a member of the most reliably graded class of question types. Not reliably graded are questions that require graphical representations, where students drawing lines into graphs performs worse-than-random. It should be noted that graphing-problems involved a background grid, which is visually distracting for AI-vision as all details of the graphs are encoded regardless of them being foreground or background.
@@ 4.3 Filtering based on partial-credit threshold
Table 2:  Quality measures for the AI grading results when employing minimum partial-credit thresholds.
Threshold Accepted Accuracy Precision Recall F1
10% 59.4% 0.84 0.92 0.82 0.87
20% 57.5% 0.85 0.91 0.83 0.87
30% 51.6% 0.84 0.89 0.81 0.85
40% 47.2% 0.82 0.87 0.77 0.82
50% 46.8% 0.82 0.87 0.77 0.82
60% 36.1% 0.83 0.82 0.73 0.77
70% 32.9% 0.84 0.82 0.73 0.77
80% 28.5% 0.86 0.80 0.72 0.76
90% 27.5% 0.86 0.80 0.73 0.76
100% 27.4% 0.86 0.80 0.73 0.76
When we screen AI-assigned partial credits by raising the minimum fraction of points we’ll trust from 10% up to full credit, we see a classic precision-recall tradeoff, see Table 2. At low cutoffs (e.g. accepting anything above 10%), the AI “covers” the most items, but some of those are borderline or incorrect guesses. So, while recall is high (we capture nearly all true positives) and overall accuracy is solid, precision drops a bit because more false positives slip through. As we demand higher partial credit (50% or full marks), precision climbs (nearly every accepted item is indeed correct, based on the ground truth), but recall falls, and we hand off more work to human graders. The F1-score (the harmonic mean of precision and recall) peaks at moderate thresholds, suggesting that a middle ground balances catching most true solves without too many over-promises.
@@
Fig.  12:  Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the full exam. The panels, show the results when only accepting problem parts for which the AI assigned at least $10\%$ partial credit, at least half credit, and full credit, respectively.
@@
The linear regressions of total AI-points versus true points echo this pattern, see Fig. 12. With a generous 10% cutoff, the slope is shallower and the intercept higher, reflecting consistent under?scoring at low thresholds and some baseline offset, yet the model still explains about 95% of the variance in true class totals. At 50% partial credit, the slope moves closer to unity, intercept shrinks, and the coefficient of determination ($R^{2}$) rises modestly, meaning the AI’s total tallies align more faithfully with human scores when we only count mid- to high-confidence judgments. Finally, when we accept only full-credit items, the fit is almost perfect (slope $\approx 1$, intercept minimal, $R^{2}\approx 0.98$), but we’re covering less than a third of all parts — so the automated grading chunk is small but highly reliable.
@@
Table 3 shows the effect of a 50&-partial-credit “mostly correct” filter on the multiclass quality metrics for different question types. The improvements notable in the linear regression are not reflected in these measures. Instead, long answers appear to be graded much worse than before, however, looking at the small decrease in number of samples, it turns out that most of these answers had been mostly correct to begin with.
@@ 4.4 Filtering based on Bayesian statistics
Fig.  13:  The matrix $\text{Risk}_{ij}$; differences are indicated by colors from blue (perfect agreement) to red (opposite of expectations). The dark red indicates the cases in which the AI assigned more than the maximum points available.
@@
Fig.  14:  Scatterplot and linear regression of the AI-assigned versus the TA-assigned points for the full exam, filtered for problem parts where the AI-assigned partial credit deviates no more than 50% from the expected partial credit.
@@
Table 4:  Quality measures for the AI grading results in order of decreasing F1, filtered for problem parts where the AI-assigned partial credit deviates no more than 50% from the expected partial credit.
@@ 4.4.1 Filtering by problem type
Probably the most straightforward way of filtering the grading decisions is to a-priory leave the grading of problem parts that include drawing or graphing to humans. Figure 15 shows the results when applying only this filter, when in addition only accepting problem parts for which the AI assigned at least $50\%$ partial credit, and when in addition only accepting less than 50% risk based on Bayesian statistics.
Fig.  15:  Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for only textual problems. The panels, show the results when applying only this filter, when in addition only accepting problem parts for which the AI assigned at least $50\%$ partial credit, and when in addition only accepting less than 50% risk based on Bayesian statistics.
Adding filters on top of the problem-type selection leads to marginal improvements, where however the risk-based filter lowers the false positive (as evidenced by the lower offset) while only rejecting an additional 8%.
@@ 5 Discussion
The results of this study underscore both the potential and current limitations of employing artificial intelligence (AI) systems for grading and assessment in chemistry education. Consistent with prior research, we found that AI-based scoring aligns closely with human grading for textual answers and chemical reaction equations, indicating that AI can effectively manage routine assessment tasks, thus allowing instructors and teaching assistants more time for nuanced instructional activities 27, 26. However, significant discrepancies remain for question formats involving graphical representations, echoing earlier findings that even advanced multimodal AI models struggle with diagrammatic reasoning and chemical structure interpretation 47, 46.
Our page-by-page rubric-based grading approach proved practical and intuitive, providing a clear method for instructors to integrate AI into existing grading workflows. This aligns with literature emphasizing intuitive and transparent grading practices to facilitate adoption among instructors 48. Given these results, we recommend initially restricting AI-assisted grading primarily to textual responses and chemical equations, while reserving graphical and complex problem-solving tasks for human graders.
@@
  • The risk-based filter, employing Bayesian statistical methods, demonstrated greater efficiency in aligning scores and reducing TA workload. This approach aligns with recent studies advocating Bayesian and probabilistic methods in educational assessment to systematically manage grading uncertainties 39, 49. However, its complexity and lower transparency may pose challenges for student acceptance.
@@
  • The problem-type-based filter straightforwardly eliminates AI grading for historically challenging question formats, significantly improving reliability. Incorporating a risk-based filter on top provided minor additional improvements in grading alignment at a modest increase in human workload, supporting recommendations for targeted AI application in chemistry education 47.
@@
The variability observed in AI and human grader agreement highlights the need for continued oversight, particularly in high-stakes assessments. The high $R^{2}$ values observed confirm that AI grading decisions are systematic and consistent with expert judgments, supporting literature indicating that AI can reliably mirror human grading patterns under controlled conditions 26, 50.
@@
Implementing AI in formative and summative assessments also raises broader considerations about feedback quality and student perceptions. Future research must investigate student trust and perceived fairness, crucial elements influencing educational outcomes and acceptance of AI technologies. Recent studies suggest that clear explanations and transparency significantly enhance student acceptance of AI-graded assessments 50. Thus, ensuring transparency, accountability, and pedagogical effectiveness should remain central goals in AI-enhanced assessment frameworks 48, 27.
In terms of cost, grading the 296 students by TAs cost $3,500/459\times 296\approx\$2,250$. As only half of the results from the AI were deemed trustworthy, the cost using AI would be $100+\frac{1}{2}\$2,250\approx\$1,225$. Thus, AI grading would not necessarily be a huge cost-savings; instead, it would allow to continue offering open-ended exam questions even in times of rising student numbers and stagnant resources.

### v2

@@ 3.5.1 Partial-credit threshold
Experiences from earlier studies suggest that AI is more likely to grade correct answers as incorrect (i.e., produce false negatives) that the other way around (false positives) kortemeyer2025assessing. One possible filter is thus to binarize the partial grading decisions and treat negatives as suspicious. For the purpose of filtering the AI grading decisions, we turn the partial-credit judgements into binary decisions based on a threshold $t$ for “correctness.” If a result is “incorrect" or “incorrect enough” based on our threshold, we discard the AI-judgement and, in a production setting, would delegate the grading of that problem part for that student to a human grader.
@@
Cases where the AI marks a part as “correct”’ but the TA’s score falls below $t$.
@@ 3.5.2 Risk threshold
Psychometrics originated in the early 20th century as a means to standardize the measurement of psychological traits such as intelligence and aptitude. The field introduced techniques like factor analysis to identify latent variables underlying observed responses, reducing them to common dimensions. A prominent early example was the general intelligence factor, $g$, proposed to capture overall cognitive ability hart1912general, thurstone1934vectors, spearman1961general. Over time, such broad constructs gave way to domain-specific abilities shaped by learning, and the focus of psychometrics gradually shifted from evaluating individuals to evaluating the assessments themselves, providing tools to ensure the validity, reliability, and fairness of tests.
@@
where $\theta_{i}$ is the latent “ability” of student $i$, and $a_{j}$ and $b_{j}$ are the estimated discrimination and difficulty of part $j$, respectively rasch, kortemeyer2019quick.
The AI’s actual normalized score $s_{ij}$ is then compared to this expectation sinharay2006posterior, and we define the “risk” of accepting that AI judgment as the absolute deviation
@@ 3.5.3 Problem types
Previous experiences have shown that multimodal AI does often succeed in reading handwriting, but still fails to interpret scientific diagrams and graphs polverini24, kortemeyer2024grading, kortemeyer2025assessing, kortemeyer2025multilingual. A straightforward filtering mechanism is thus to exclude any problem parts that have graphical components, namely “Drawing” and “Graphing” (see Fig. 3), i.e., filter for the remaining problem parts that are purely textual or require nothing more than recognizing the position of a cross in multiple-choice problems..
@@ 4 Results
4.1 Inter-run reliability of points
On total scores per student, the AI showed high run-to-run consistency: ICC(A,1)=0.967=0.967 for a single run and ICC(A,5)=0.993=0.993 if one were to average five runs. The within-target repeatability SD was $S_{w}=1.92$ points, implying a 95% repeatability coefficient $\mathrm{RC}=5.33$ points; i.e., for two independent runs, 95% of absolute differences in a given student’s total would be at or below $5.33$ points on the raw scale. Note that the 95% repeatability coefficient is a conservative bound; under the same model the typical absolute difference between two independent runs is about $2.17$ points ($\approx 3.6\%$ of $60$), and averaging five runs would reduce the 95% bound to $2.39$ points ($\approx 4.0\%$). Rank concordance of totals across runs was likewise high (Kendall’s $W=0.959$), and level agreement was strong (pairwise Lin’s CCC on totals: median $0.977$, IQR $0.045$), indicating both similar ordering and minimal mean/scale bias between runs. As an intuitive dispersion summary, the per-student SD of totals across runs had a median of $1.54$ points (IQR $1.13$).
[…]
@@ 4.2 Qualitative results
A comparison of the exam sheets graded by the TAs and score sheets provided by AI, revealed some interesting features of our AI-grading mechanism. This are anecdotal in nature and possibly specific to our grading mechanism (see Sect. 6).
@@ 4.2.2 False negatives
Fig. 6: An example of a false negative (problem 6-C-b).
@@
Fig. 7: An example of a false negative, where AI assigned only partial credit (problem 2-C-a).
@@
Fig. 8: Incorrect interpretation of a multiple-response problem (problem 6-A).
In Fig. 8, AI does not correctly recognize the position of a cross. It claims that the second-to-last answer contains the cross in the left column named “richtig” (correct) and deducts points. It is unclear how this may have happened, except that the student markings are not really crosses.
Fig. 9: Incorrect interpretation of labels (problem 4-A-a).
@@ 4.3 Unfiltered quantitative results
Fig. 10: Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the six problems on the exam. Each data point represents one student.
As Fig. 10 shows, for the individual six problems, the AI-derived scores correlated with human scores at coefficients of determination ($R^{2}$) ranging from $0.61$ to $0.85$. The reliability on a per-problem basis is low and not suitable for grading without additional measures, yet the aggregate total score shown in the top-left panel of Fig. 11 achieved an $R^{2}$ of $0.91$ — that this $R^{2}$ is “bigger than the sum of its parts” may initially appear surprising. The enhancement results from the partial cancellation of random scoring errors across problems: when individual deviations offset one another in the sum, overall noise diminishes relative to the true-score variance. The resulting increase in the signal-to-noise ratio produces a more reliable composite measure of student proficiency than any single item alone. In this context, by “noise” we mean unsystematic item-level misgrading — some items scored too high and others too low — so aggregation partially cancels these opposite errors, increasing the reliability of the total (as predicted by the Spearman–Brown prophecy in Classical Test Theory). This argument, borne out by our result, assumes that item-level errors are approximately mean-zero and only weakly correlated across problems; systematic biases would not cancel.
Fig. 11: Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the full exam. The maximum achieved score was 52; as typical for exams in the German-speaking university tradition, no student achieved full points.
As Table 1 shows, the multiclass quality of the grading varies between question types. The first column lists the question type with additional rows across all question types and a row for combined question types. The second column shows the number of these parts, the third column the number of student-part samples, the fourth column the number of classes $n$, the four following columns the quality measures, and the rightmost column the normed F1 (Equation 1). The table is sorted in order of decreasing normed F1 to highlight the question types that are most accurately graded.
Overall, these quality measures might appear low, however, one needs to keep in mind that these are multiclass measures. If a problem part is worth two points, and possible points are 0.0, 0.25, …, 2.0, these are nine classes (“notches”), and a true positive would only be counted if the fractional credit agrees exactly; if the ground truth were 1.25 points and the AI assigned 1.5 points, the AI-grading would be considered just as incorrect as if it had assigned 0.25 points.
Table 1: Multiclass quality measures for the AI grading results in order of decreasing F1.
@@
Not surprisingly for a language model, short and long textual answers emerge as the most reliably graded one. Fortunately for chemistry education, equations of chemical reactions are also a member of the most reliably graded class of question types. Not reliably graded are questions that require graphical representations, where students are drawing lines into graphs; here, the system performs worse-than-random. It should be noted that these graphing problems involved a background grid, which is visually distracting for AI-vision as all details of the graphs are encoded regardless of them being foreground or background.
@@ 4.4 Filtering based on partial-credit threshold
Table 2 reports the performance of selective automation for AI-assigned partial credit as a function of the acceptance threshold $t$ For each student-part cell, the AI proposes a fractional score; if that score is at least $t$, the item is auto-accepted, otherwise it is deferred to human grading. “Accepted” gives the resulting coverage (percentage of items handled by the AI). “MAE” is the mean absolute error in points computed only over accepted items. “Under” and “Over” are the percentages of accepted items on which the AI under- or over-credits relative to human grading. “Under > 0.25 Pts” flags accepted items that are off by what is generally more than one grading notch (0.25 points) on the under-credit side. “Bias Pts” is the mean difference (AI - ground truth, in points) on the accepted set, with negative values indicating a tendency to short students and positive values indicating generosity.
Table 2: Accepted = share of items auto-graded (AI score $\geq t$); MAE = mean absolute point error on accepted items; Under/Over = share of accepted items where AI is below/above the human score; Under > 0.25 Pts = share of accepted items under-credited by more than 0.25 points; Bias Pts = mean(AI - ground truth) on accepted items.
Threshold $t$ Accepted MAE Pts Under Under > 0.25 Pts Over Bias Pts
0 100% 0.21 24% 15% 14% -0.08
0.1 59% 0.22 21% 12% 24% 0
0.2 57% 0.21 20% 12% 24% 0.01
0.3 52% 0.21 18% 10% 24% 0.04
0.4 47% 0.18 11% 6% 26% 0.08
0.5 47% 0.18 11% 6% 25% 0.08
0.6 36% 0.16 6% 3% 26% 0.11
0.7 33% 0.14 4% 1% 23% 0.12
0.8 28% 0.13 1% 0% 22% 0.12
0.9 28% 0.12 0% 0% 21% 0.12
1 27% 0.12 0% 0% 20% 0.12
A threshold $t=0$ means unfiltered. As the threshold increases, the system becomes more conservative: coverage falls, but the accepted items are cleaner. The MAE is always smaller than a grading notch, and it improves further with increasing threshold. The “Under” rate, alongside with the “Under > 0.25 Pts” rate contract to zero at high thresholds; in other words, tightening $t$ effectively suppresses the student-harm modes, so both the frequency and the magnitude of under-credit among the auto-graded items decrease. The bias on the accepted set shifts from slightly negative at very low thresholds toward small positive values at higher thresholds, indicating that once we keep only higher-scored cases, the AI is no longer prone to short students on average.
From a grade-integrity and workflow perspective, the residual “Over” rate on the accepted set remains a salient consideration: even when student-harm is low, some accepted items may be graded a bit generously. The table therefore supports policy choices by making the trade-off explicit: one can select the highest threshold that keeps “Under > 0.25 Pts” below a tolerance deemed acceptable (e.g., 1-5%) while still delivering useful coverage. In practice, this yields a defensible operating point that protects students against under-credit on auto-graded work, limits systematic bias, and preserves human review for the low-score and borderline cases where it adds the most value.
@@
Fig. 12: Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for the full exam. The panels, show the results when only accepting problem parts for which the AI assigned at least $10\%$ partial credit, at least half credit, and full credit, respectively.
@@
The linear regressions of total AI-points versus true points echo this pattern, see Fig. 12. With a generous 10% cutoff, the slope is shallower and the intercept higher, reflecting consistent under-scoring at low thresholds and some baseline offset, yet the model still explains about 95% of the variance in true class totals. At 50% partial credit, the slope moves closer to unity, intercept shrinks, and the coefficient of determination ($R^{2}$) rises modestly, meaning the AI’s total tallies align more faithfully with human scores when we only count mid- to high-confidence judgments. Finally, when we accept only full-credit items, the fit is almost perfect (slope $\approx 1$, intercept minimal, $R^{2}\approx 0.98$), but we’re covering less than a third of all parts — so the automated grading chunk is small but highly reliable.
@@
Table 3 shows the effect of a 50%-partial-credit “mostly correct” filter on the multiclass quality metrics for different question types. The improvements notable in the linear regression are not reflected in these measures. Instead, long answers appear to be graded much worse than before, however, looking at the small decrease in number of samples, it turns out that most of these answers had been mostly correct to begin with.
@@ 4.5 Filtering based on Bayesian statistics
Fig. 13: The matrix $\text{Risk}_{ij}$; differences are indicated by colors from blue (perfect agreement) to red (opposite of expectations). The dark red indicates the cases in which the absolute difference between the AI-assigned and the expected score was bigger than one, which can happen if the AI assigned more than the maximum points available or if it assigned negative points (see Sect. 4.2.2); the former kind of overflow happened twice for problem part 2-B and once for problem part 3-B-a (both numerical problems), the latter once for problem part 5-C-a (a multiple-choice problem)).
@@
Fig. 14: Scatterplot and linear regression of the AI-assigned versus the TA-assigned points for the full exam, filtered for problem parts where the AI-assigned partial credit deviates no more than 50% from the expected partial credit.
@@
Table 4: Multiclass quality measures for the AI grading results in order of decreasing F1, filtered for problem parts where the AI-assigned partial credit deviates no more than 50% from the expected partial credit.
@@ 4.5.1 Filtering by problem type
Probably the most straightforward way of filtering the grading decisions is to a priory leave the grading of problem parts that include drawing or graphing to humans. Figure 15 shows the results when applying only this filter, when in addition only accepting problem parts for which the AI assigned at least $50\%$ partial credit, and when in addition only accepting less than 50% risk based on Bayesian statistics.
Fig. 15: Scatterplots and linear regressions of the AI-assigned versus the TA-assigned points for only textual problems. The panels show the results when applying only this filter, when in addition only accepting problem parts for which the AI assigned at least $50\%$ partial credit, and when in addition only accepting less than 50% risk based on Bayesian statistics.
Adding filters on top of the problem-type selection leads to marginal improvements, where however the risk-based filter lowers the false positives (as evidenced by the lower offset) while only rejecting an additional 8%.
@@ 5 Discussion
The results of this study underscore both the potential and current limitations of employing artificial intelligence (AI) systems for grading and assessment in chemistry education. Consistent with prior research, we found that AI-based scoring aligns closely with human grading for textual answers and chemical reaction equations, indicating that AI can effectively manage routine assessment tasks, thus allowing instructors and teaching assistants more time for nuanced instructional activities Fernandez2024, Floden2025. However, significant discrepancies remain for question formats involving graphical representations, echoing earlier findings that even advanced multimodal AI models struggle with diagrammatic reasoning and chemical structure interpretation Ryan2019, kortemeyer2025multilingual.
Our page-by-page rubric-based grading approach proved practical and intuitive, providing a clear method for instructors to integrate AI into existing grading workflows. This aligns with literature emphasizing intuitive and transparent grading practices to facilitate adoption among instructors Azad2020. Given these results, we recommend initially restricting AI-assisted grading primarily to textual responses and chemical equations, while reserving graphical and complex problem-solving tasks for human graders.
@@
  • The risk-based filter, employing Bayesian statistical methods, demonstrated greater efficiency in aligning scores and reducing TA workload. This approach aligns with recent studies advocating Bayesian and probabilistic methods in educational assessment to systematically manage grading uncertainties kortemeyer2025assessing, Zitzmann2024. However, its complexity and lower transparency may pose challenges for student acceptance.
@@
  • The problem-type-based filter straightforwardly eliminates AI grading for historically challenging question formats, significantly improving reliability. Incorporating a risk-based filter on top provided minor additional improvements in grading alignment at a modest increase in human workload, supporting recommendations for targeted AI application in chemistry education Ryan2019.
@@
The variability observed in AI and human grader agreement highlights the need for continued oversight, particularly in high-stakes assessments. The high $R^{2}$ values observed confirm that AI grading decisions are systematic and consistent with expert judgments, supporting literature indicating that AI can reliably mirror human grading patterns under controlled conditions Floden2025, Emirtekin2025. Our inter-run reliability analysis indicates that averaging several, independent grading runs could further improve the accuracy of the results.
@@
Implementing AI in formative and summative assessments also raises broader considerations about feedback quality and student perceptions. Future research must investigate student trust and perceived fairness, crucial elements influencing educational outcomes and acceptance of AI technologies. Recent studies suggest that clear explanations and transparency significantly enhance student acceptance of AI-graded assessments Emirtekin2025. Thus, ensuring transparency, accountability, and pedagogical effectiveness should remain central goals in AI-enhanced assessment frameworks Azad2020, Fernandez2024.
The purpose of our study was to explore the feasibility of offering meaningful, open-ended assessment more frequently to rising numbers of students. Thus, the goal is efficiency, not necessarily cost savings. Nevertheless, it is instructive to look at cost: grading the 296 students by TAs cost $\mathdollar 3,500/459\times 296\approx\mathdollar 2,250$. As only half of the results from the AI were deemed trustworthy, the cost using AI would be $\mathdollar 100+\frac{1}{2}\mathdollar 2,250\approx\mathdollar 1,225$. Not counted is the time required for depersonalization, as that would not be necessary outside of our IRB-protocol, and the time required for scanning (as this was done by senior personnel for our study, it cost \$1,400 in hourly-salary equivalent; it should be noted, though, that already today, many departments routinely scan their exams for evidence or for paperless grading). In both scenarios, students may appeal their scores, which requires additional faculty and staff time; this might happen more frequently with AI-assigned scores. Thus, AI grading would not necessarily be a huge cost-savings; instead, it would allow instructors to continue offering open-ended exam questions even in times of rising student numbers and stagnant resources.
6 Limitations
We did not perform parallel double scoring to estimate inter-rater reliability. Implementing a second, fully independent scoring stream would have required suppressing the calibration and adjudication steps described above and preventing the instructor from recalling prior decisions, thereby deviating from the authentic grading workflow. Consequently, our labels represent the as-implemented institutional scoring rather than an independently replicated consensus. Residual rater error is possible; however, the design (single rater per item, upfront co-grading with a detailed rubric, documented adjudications of edge cases, instructor audit, and a post-exam appeals process) was intended to minimize idiosyncratic variance. Future work could assess inter-rater reliability on a stratified subsample or via a generalizability analysis, but this was outside the scope of the present, practice-grounded study.
The results presented here are specific to our exam, the model used (GPT-o4-mini), and the prompting; while we have found similar results with GPT-4o grading a thermodynamics exam kortemeyer2025assessing and GPT-5 grading a calculus exam kortemeyer2025artificial, results are not necessarily generalizable across models and prompts. For example, the negative points assigned to multiple-choice problems discussed in Sect. 4.2.2 could likely have been avoided with a more carefully formulated rubric (“prompting”), however, post-hoc manipulation of the rubric might defeat the purpose of modeling an authentic scenario.